In an integrated science class of 76 students, 8 are chemistry-education; 3 physics-education; 39 biology-education; and 26 mathematics-education. What is the probability of drawing a:
a. Biology-education students from the pool?
b. Chemistry or physics-education students from the pool?
c. Mathematics or Physics-education student from the pool?

Answer by jorel555(1290)

a)39/76=.5132 probability
b)(8+3)/76=.145 probability
c)(26+3)/76=.382 probability
☺☺☺☺